According to the University of Nevada Center for Logistics Management, 8% of all merchandise sold in the United States gets returned. A Houston department store sampled 60 items sold in January and found that 11 of the items were returned a. Construct a point estimate of the proportion of items returned for the population of sales transactions at the Houston store 1833 (to 4 decimals) b. Construct a 99% confidence interval for the proportion of returns at the Houston store. ) (to 4 decimals) c. Is the proportion of returns at the Houston store significantly different from the returns for the nation as a whole? Provide statistical support for your answer 0.08 0.08 z-value 3 (to 2 decimals) p-value (to 4 decimals) Reject the null. No, we have enough evidence to prove that Houston is different from the national average
Answer #1

Please don't hesitate to give a "thumbs up " for the answer, in case you're satisfied with it.

a.

p = .08
n = 60
x = 11

a. The point estimate is therefore, x/n = 11/60 = .1833

b. 99% CI is given by : p +/- Z*sqrt(p*p'/n) = 0.1833 +/- 2.92*sqrt(.1833*(1-.1833)/60) = -0.0659 to 0.2259

[since the population is finite]

c.

Ho: p0 = .08
Ha: p0!= .08

z-value is : (p^-p)/sqrt(p*p'/n) = (.1833-.08)/sqrt(.08*.92/60) = 2.95

p-value is 0.0032

(we have p-value of .0032<< .05, hence we reject the null hypothesis)

Reject the null. No, we have enough evidence....